OpenStudy (anonymous):

Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Laura has the movie

jigglypuff314 (jigglypuff314):

Welcome to OpenStudy! :) rate = slope = m = 0.5 constant = 2 therefore when using the slope intercept form: y = mx + b you would plug in m = 0.5 and b = 2 to get your equation :)

OpenStudy (anonymous):

thank you :). and will it be c(x) = 2.50x?

jigglypuff314 (jigglypuff314):

not quite, because numbers with variables attached to them cannot be added to numbers without variables attached to them so c(x) = 0.5x + 2

OpenStudy (anonymous):

c(x) = 2.00x + 0.50?

jigglypuff314 (jigglypuff314):

no it would just be c(x) = 0.5x + 2 because y = mx + b then plug in m = 0.5 and b = 2 to get your equation

jigglypuff314 (jigglypuff314):

0.5 would be the one attached to the variable because it changes based on how many nights she keeps the movie and 2 would not be with a variable because it would not be changing if that helps...

OpenStudy (anonymous):

but those are some of my options

jigglypuff314 (jigglypuff314):

hmm.... what are all your options then? :/

OpenStudy (anonymous):

c(x) = 2.00x + 0.50 c(x) = 2.00 + 0.50x c(x) = 2.50x c(x) = (2.00 + 0.50)x

jigglypuff314 (jigglypuff314):

ah yes, okay, by commutative property of addition c(x) = 0.5x + 2 can also equal c(x) = 2 + 0.5x

jigglypuff314 (jigglypuff314):

commutative property of addition is like when 2+1 = 1+2

OpenStudy (anonymous):

so c?

jigglypuff314 (jigglypuff314):

c(x) = 0.5x + 2 can also equal c(x) = 2 + 0.5x so c(x) = 2.00 + 0.50x

OpenStudy (anonymous):

omg thanx jig can uu hep with another 1?

jigglypuff314 (jigglypuff314):

I kinda have to go to school now >,< sorry

OpenStudy (anonymous):

aww man ok